# TLControl: Trajectory and Language Control for Human Motion Synthesis

### 28 Nov 2023 v1

@@ 3.1 Part-based VQ-VAE
Learning discrete representations has proven effective in motion synthesis tasks in recent works [14, 57]. These methods use VQ-VAE [48] for discretizing a continuous motion space, facilitating the capturing of complex motion patterns in a structured manner and enhancing the generation of coherent and high-fidelity motion sequences. However, previous works [14, 57] encode the full body into the discrete motion space, treating the human body as a holistic entity. Instead, in this paper, we propose to learn a better representation of part-based motion priors, which have demonstrated impressive results for computer character animation [43, 4, 18]. To do so, we propose a part-based VQ-VAE that disaggregates the motion at the human body-part level for learning a well-structured and compact latent space.
@@ 3.2 Masked Trajectory Transformer
We train the MTT encoder $\bm{\mathcal{T}}$ to encode user-provided observations into a compressed latent space $\mathbf{Z}=\bm{\mathcal{T}}(\mathbf{R}^{\prime},\bm{L})$. Then, we employ a decoders on $\mathbf{Z}$, which focuses on inferring a list of code indices $\hat{I}$ in the codebooks from $\mathbf{Z}$, where the codes of these indices should form the quantized latent codes $\hat{\mathbf{Q}}_{0}$ that can be decoded to corresponding motion sequences through our decoder. The loss function for training the MTT is:
$\mathcal{L}=\mathbb{E}_{\hat{I}\sim P(I)}[-\text{log}P(\hat{I}|R^{\prime},L)]+||\mathcal{\mathbf{D}}(\hat{\mathbf{Q}}_{0})-\mathbf{J}||_{2}$ (3)
@@ 4 Experiment
In the part-based VQ-VAE, the settings of the encoder and the decoder module are based on T2M-GPT [57] with the downsampling $s$ factor set to 4. For the masked trajectory transformer, we use a frozen CLIP-ViT-B/32 [36] model for pre-processing the text prompt $\bm{L}$, and then a 4-layer transformer processes the text information and control trajectories $\bm{R^{\prime}}$ into the latent $\bm{Z}$, followed by a 3-layer transformer decodes $\bm{Z}$ to the codes indices. The embedding dimension of these two stage transformers are 512 and 256 respectively. The networks are trained using AdamW optimizer with learning rate decayes from 1e-4 to 1e-7 with a batch size of 64. For the run-time optimization step, we use Limited-Memory BFGS [26] as the optimization algorithm with learning rate set to 0.1. The accuracy criteria is set to 1E-6 and we provide an ablation study of the criteria setting in  4.3.4.
@@ 4.3.1 Motion Latent Space
We investigate the learned part-based VQ-VAE for motion embedding. In Fig. 5, we randomly choose 10 test cases from Humanml3D test set, and process them with MMT for achieving the corresponding features. We present the result of using t-SNE [49] for visualizing these features, offering insights into the structure of the acquired motion latent space. Our part-based VQ-VAEcontributes a compact motion latent space while preserving the semantics described by language inputs, as evidenced in Fig. 5.
@@ 4.3.2 Trajectory Incompleteness
Figure 8: Running time statistics of our optimization when applying different controlled joints.
We investigate the performance of our method under different levels of trajectory incompleteness, for which we simulate various incompleteness scenarios by adjusting the ratio of random masking. We include the reconstructed trajectories during the optimization stage. As revealed in Fig. 7, our model’s performance degrades gracefully even as the control trajectories become more incomplete. As a reference, note that even with full joint trajectories, the performance of previous methods [53] in terms of FID and R-precision are $2.614$ and $0.606$.
@@ 4 Implementation Details
In implementing the masked trajectory transformer, we employ a frozen CLIP-ViT-B/32 [36] model for the initial processing of the text prompt, which translates it into a 512-dimensional vector representative of language features. The joint trajectories are grouped with every 4 closest waypoints being bundled into a single temporal token. This bundling is reflective of our choice to use a downsampling rate of 4 within the part-based VQ-VAE framework. Each waypoint group is then mapped onto a 512-dimensional feature space. A standard 4-layer transformer encoder is used to integrate the language and trajectory features into a cohesive latent space, where tokens comprise both the language feature vector and the feature representations of the trajectory at each temporal token. Following this, a 3-layer transformer encoder is designed to transform the latent space into a sequence of logits, which are used to determine the code indices. We then apply a softmax function followed by an argmax operation to select the most probable code indices, which serve as the initial full-body state for the VQ-decoder. For the masking strategies, each iteration within a batch has a 50% probability of using continuous trajectory masking or joint-level masking. The proportion of continuous trajectory masking is gradually increased, starting at 0% and steadily advancing to 75%.

### 12 Dec 2023 v3

@@ 3.1 Part-based VQ-VAE
Learning discrete representations has proven effective in motion synthesis tasks in recent works [14, 58]. These methods use VQ-VAE [49] for discretizing a continuous motion space, facilitating the capturing of complex motion patterns in a structured manner and enhancing the generation of coherent and high-fidelity motion sequences. However, previous works [14, 58] encode the full body into the discrete motion space, treating the human body as a holistic entity. Instead, in this paper, we propose to learn a better representation of part-based motion priors, which have demonstrated impressive results for computer character animation [44, 4, 18, 34]. To do so, we propose a part-based VQ-VAE that disaggregates the motion at the human body-part level for learning a well-structured and compact latent space.
@@ 3.2 Masked Trajectory Transformer
We train the MTT to process the user-provided observations $(\mathbf{R}^{\prime},\bm{L})$ to infer quantized latent VQ-VAE motion codes corresponding to the ground truth motions $\mathbf{J}$. Specifically, the MTT infers indices $I$. Indexing with $I$ into the part codebooks $\{C_{k}\}$ produces the concatenated quantized latent $\hat{\mathbf{Q}}_{0}$, which may be decoded using the frozen VQ-VAE decoder $\bm{\mathcal{D}}$. Thus, the loss function for training the MTT is:
$\mathcal{L}=\mathbb{E}_{\hat{I}\sim P(I)}[-\text{log}P(\hat{I}|R^{\prime},L)]+||\bm{\mathcal{D}}(\hat{\mathbf{Q}}_{0})-\mathbf{J}||_{2}.$ (3)
Please refer to our Suppl. Materials for more details.
@@ 4 Experiment
In the part-based VQ-VAE, the settings of the encoder and the decoder module are based on T2M-GPT [58] with the downsampling $s$ factor set to 4. For the masked trajectory transformer, we use a frozen CLIP-ViT-B/32 [37] model for pre-processing the text prompt $\bm{L}$, and then a 4-layer transformer processes the text information and control trajectories $\bm{R^{\prime}}$ into the latent features, followed by a 3-layer transformer decodes the latent features into the codes indices. The embedding dimension of these two stage transformers are 512 and 256 respectively. The networks are trained using AdamW optimizer with learning rate decayes from 1e-4 to 1e-7 with a batch size of 64. For the run-time optimization step, we use Limited-Memory BFGS [26] as the optimization algorithm with learning rate set to 0.1. The accuracy criteria is set to 1E-6 and we provide an ablation study of the criteria setting in  4.3.4.
@@ 4.3.1 Motion Latent Space
We investigate the learned part-based VQ-VAE for motion embedding. In Fig. 5, we randomly choose 10 test cases from Humanml3D test set, and process them with MTT for achieving the corresponding features. We present the result of using t-SNE [50] for visualizing these features, offering insights into the structure of the acquired motion latent space. Our part-based VQ-VAEcontributes a compact motion latent space while preserving the semantics described by language inputs, as evidenced in Fig. 5.
@@ 4.3.2 Trajectory Incompleteness
Figure 8: Running time statistics of our optimization when applying different controlled joints. “Upper Body” includes the joints of the hands and the head joint. “Lower Body” includes the joints of two feet and the joint of the pelvis.
We investigate the performance of our method under different levels of trajectory incompleteness, for which we simulate various incompleteness scenarios by adjusting the ratio of random masking. As revealed in Fig. 7, our model’s performance degrades gracefully even as the control trajectories become more incomplete. As a reference, note that even with full joint trajectories, the performance of previous methods [54] in terms of FID and R-precision are $2.614$ and $0.606$.
@@ 5 Conclusion
6 Acknowledgements
The authors would like to thank Yiming Xie for preparing sample cases in the comparisons. Dinesh Jayaraman was supported by NSF CAREER Award 2239301.
@@ 4 Implementation Details
In implementing the masked trajectory transformer, we employ a frozen CLIP-ViT-B/32 [37] model for the initial processing of the text prompt, which translates it into a 512-dimensional vector representative of language features. The joint trajectories are grouped with every 4 closest waypoints being bundled into a single temporal token. This bundling is reflective of our choice to use a downsampling rate of 4 within the part-based VQ-VAE framework. Each waypoint group is then mapped onto a 512-dimensional feature space. A standard 4-layer transformer encoder is used to integrate the language and trajectory features into a cohesive latent space, where tokens comprise both the language feature vector and the feature representations of the trajectory at each temporal token. Following this, a 3-layer transformer encoder is designed to transform the latent space into a sequence of logits for determining the code indices. To transform logits into selected code indices, we follow the common idea of applying softmax and then argmax, but in a differentiable manner: we use the Gumbel-Softmax [^1] Gumbel-Softmax:  https://pytorch.org/docs/stable/generated/torch.nn.functional.gumbel_softmax.html for producing a one-hot decision vector $\mathbf{v}$, followed by selecting the most probable code index using $\mathbf{C}_{k}\cdot\mathbf{v}$ where $\mathbf{C}_{k}\in\mathbb{R}^{d\times|\mathbf{C}|}$ is the codebook for a joint group containing $|\mathbf{C}|$ distinct codes, each of dimension $d$. Then the selected codes are concatenated to serve as the initial full-body state for the VQ-decoder.
For the masking strategies, each iteration within a batch has a 50% probability of using continuous trajectory masking or joint-level masking. The proportion of continuous trajectory masking is gradually increased, starting at 0% and steadily advancing to 75%.